# Who is Responsible? Data, Models, or Regulations, A Comprehensive Survey on Responsible Generative AI for a Sustainable Future

### v6

@@ Abstract.
Generative AI is rapidly moving from research to deployment, elevating the need for responsible development, evaluation, and governance. We conduct a PRISMA-guided review of 232 studies (November 2022–December 2025), spanning large language models, vision-language models, diffusion models, and agentic pipelines. We make four contributions: (1) the first survey bridging governance principles, technical evaluation, and domain deployment across all four system types; (2) a ten-criterion rubric (C1–C10) scoring major AI safety benchmarks on risk-surface coverage, paired with a policy crosswalk mapping benchmarks to regulatory requirements; (3) twelve lifecycle KPIs, explainability guidance for foundation models, and a testbed catalogue; and (4) domain-specific analysis across healthcare, finance, education, arts, agriculture, and defense. Three findings emerge: benchmark coverage is dense for bias and toxicity but sparse for privacy, provenance, deepfakes, and system-level failures in agentic settings; evaluations remain largely static and task-local, limiting audit portability; and inconsistent documentation complicates cross-release comparison. We outline a research agenda prioritizing adaptive multimodal evaluation, privacy and provenance testing, deepfake risk assessment, calibration reporting, versioned artifacts, and continuous monitoring. This survey offers a structured path to align generative AI evaluation with governance needs for safe and accountable deployment.
Project page: Project page
@@ 1. Introduction
Generative AI (GenAI) systems, including large language models (LLMs), vision language models (VLMs), diffusion models, and emerging agentic pipelines, are rapidly transitioning from research prototypes to production deployments that generate text, images, code, and autonomous actions across healthcare, finance, education, and other sectors (Hadi et al., 2023; Hajkowicz et al., 2023). This accelerated deployment heightens the need for responsible development practices that address GenAI-specific risks, including hallucination and unverifiability, prompt injection and jailbreak vulnerabilities, data memorization and leakage, deepfake and media integrity threats, and system-level failures in tool-augmented settings.
@@
Responsible AI (RAI) (Stryker, 2024) has emerged as a central area of research and practice to address these challenges. RAI frameworks emphasize fairness, bias mitigation, privacy protection, security, and the safeguarding of human rights (Peters et al., 2020). By embedding these principles throughout the AI lifecycle, including design, training, deployment, and governance, RAI seeks to realize the societal benefits of AI while minimizing potential harms. While these principles were initially developed for predictive systems, generative models introduce qualitatively new challenges: outputs are open-ended and difficult to enumerate, stochastic decoding produces non-deterministic behaviour, multimodal synthesis creates cross-modal failure modes, and agentic tool use extends the attack surface beyond the model itself. Despite rapid adoption, with 74% of organizations reporting the use of GenAI, only 26% report having a comprehensive organization-wide responsible AI strategy (Capgemini Research Institute, 2024). This imbalance underscores the need to align the development and deployment of these systems with ethical standards, fairness, privacy, security, and societal values.
@@
There has been substantial progress on RAI from both industry (Google AI, 2024; Stryker, 2024; Microsoft, 2024; Anthropic, 2024; Jiang et al., 2025) and academia (Lu et al., 2024b; Dignum and Confidential, 2019; Schiff et al., 2020; Huang et al., 2025). Nevertheless, notable gaps remain in alignment with human values and in operational practice. The academic literature often separates conceptual frameworks from practical implementations, as summarized in Table 1. Conceptual work provides principles and guidelines, whereas applied studies demonstrate deployment in specific contexts. The recent MIT AI Risk Repository analysis (Slattery et al., 2024) indicates that current AI safety frameworks, including those referenced in Table 5, cover only about 34% of risks identified by governance bodies such as the NIST AI Risk Management Framework (of Standards and, NIST), the EU AI Act (Outeda, 2024), and ISO/IEC 42001 (ISO/IEC, 2021b). This gap between governance coverage and practical implementation highlights the need for an integrated review that connects frameworks to implementation strategies and to evaluation methods, including explainable AI and AI-ready testbeds.
In this study, we conduct a comprehensive literature review of responsible GenAI (ResGenAI) frameworks, governance guidelines, evaluation benchmarks, and operational risk assessment methodologies. Our scope centres on the application of RAI principles to generative systems, including LLMs, VLMs, diffusion models, and agentic/RAG pipelines. While we draw on the broader RAI literature for foundational constructs (e.g., fairness, privacy, safety, transparency, accountability), our analysis targets ResGenAI specific risk mechanisms such as hallucination and unverifiability, prompt injections, jailbreaking and tool-use failures, data memorization and leakage, provenance and deepfakes or media integrity risks, intellectual property concerns, and system-level failures in GenAI applications. Despite substantial prior work, ResGenAI guidance remains fragmented across principles, processes, and tools (Table 1). Consequently, there is a pressing need to translate regulatory obligations into concrete engineering controls and to produce portable audit evidence for post-market monitoring.
@@
Figure 1. Responsible GenAI landscape organised into three pillars: Challenges (Section 3), Strategies (Section 4), and Applications (Section 5). Technical challenges span data, model/system, and emergent-threat layers; strategies encompass benchmarks, KPIs, explainability, testbeds, governance, and toolkits; applications map domain-specific risks and controls across six sectors.
Contributions. Our key contributions in this work are:
@@
    To the best of our knowledge, this is the first review article that simultaneously spans governance principles, technical evaluation, and domain deployment for GenAI. We synthesize findings across the NIST AI RMF, EU AI Act, ISO/IEC 42001, and other instruments listed in Table 6, covering LLMs, VLMs, diffusion models, and RAG topics. As shown in Table 1, no prior survey bridges all four thematic areas (CS theory, application, policy, GenAI) and all four novelty dimensions (taxonomy, scope, evaluation, methods) simultaneously.
@@
    We introduce a ten-criterion rubric (C1–C10) that scores widely used AI safety benchmarks on risk-surface coverage and assurance quality, and a policy crosswalk (Section 4.2.1) that maps each benchmark to the governance requirements it can plausibly evidence. The closest efforts (Liu et al., 2025b; Ferdaus et al., 2026) partially address evaluation taxonomy or trust assessment, but neither links benchmarks to regulatory audit-evidence needs.
@@
    We apply the RAI lens to real-world applications and domains (Section 5, demonstrating how technical advancements can be effectively aligned with governance frameworks and regulatory requirements.
@@ Methods
(Mehrabi et al., 2021a) ACM CSUR 2021 Bias and Fairness x x x x x x
@@
(He et al., 2025b) ACM CSUR 2025 LLM Agent Security x x x x
@@
(Liu et al., 2025b) arXiv 2025 LLM Safety Evaluation x x x x
@@
Column definitions CS Theory: conceptual frameworks, taxonomies, or algorithmic contributions. Application: domain-specific deployment or case studies. Policy: governance instruments, regulatory analysis, or compliance guidance. GenAI: explicit focus on generative models (LLMs, VLMs, diffusion, agentic). Taxonomy: structured risk classification beyond a flat listing of concerns. Scope: coverage of multimodal systems and agentic/tool-using pipelines. Evaluation: governance-aligned benchmarking with a scoring rubric or crosswalk. Methods: actionable engineering controls, lifecycle KPIs, or testbed guidance. ★ = partial or limited treatment.
The rest of this paper is organized as follows. Section 2 presents the literature review methodology. Section 3 provides background on ResGenAI. Section 4 reports our results, including thematic findings, benchmark assessments, key performance indicators, explainable AI, and AI-ready testbeds. Section 5 discusses ResGenAI applications, challenges, and strategies. Section 6 presents discussion and forward thoughts, and finally, Section 7 concludes with open challenges and future research directions. The taxonomy presented in Figure 1 provides a structural blueprint for this survey, establishing the thematic connections between high-level technical challenges and domain-specific applications.
@@ 2. Methodology
This review adheres to the PRISMA guidelines (Page et al., 2021) to ensure a transparent, reproducible, and rigorous synthesis of evidence. This study follows established methodological recommendations for performing systematic reviews within the computer science community (Al-Selwi et al., 2024; Boland et al., 2017a; Keele et al., 2007). To ensure comprehensive and consistent reporting, the structural framework was modelled after previous extensive surveys in the domain. The mapping procedure utilized in this review is organized into four interconnected phases that covering objectives (Section 2.1), search strategy (Section 2.2), eligibility criteria (Section 2.3), and study selection (Section 2.4) which are visualized in the methodological flow (Fig. 2) and detailed throughout Section 2. For a more exhaustive exploration of these mapping and reporting protocols, foundational works in the field (Boland et al., 2017a; Al-Selwi et al., 2024; Keele et al., 2007; Boland et al., 2017b) offer additional guidance.
Figure 2. The systematic literature mapping process utilized in this study
2.1. Research Questions Formulation
This study conducts a literature review of responsible generative AI (ResGenAI) in the post-GPT period and the synthesis of the findings. We address the following key questions:
  • RQ1: What are the dominant themes, risks, and controls in ResGenAI and multimodal systems since late 2022?
@@ 2.1. Research Questions Formulation
    Motivation: The rapid transition of GenAI from research to production has outpaced traditional risk taxonomies, necessitating a contemporary synthesis of how multimodal and agentic systems introduce qualitatively new failure modes beyond predictive ML.
  • RQ2: Which evaluation facets and metrics are most commonly reported, and where are the gaps?
@@
    Motivation: Identifying the “Evaluation Gap” is critical for determining if current safety suites adequately model the adaptive, iterative pressure applied by real-world adversaries versus merely serving as static, keyword-based refusal scripts.
  • RQ3: What evidence types (empirical, analytical, normative) underpin claims about RAI effectiveness?
@@
    Motivation: A structural deficit in evidence portability hinders institutional trust; this question seeks to evaluate if safety claims are grounded in verifiable technical artifacts or subjective, unvalidated appraisals.
  • RQ4: How can these findings be organized into a practical framework for research and governance?
@@
    Motivation: Principles without practice lead to “safetywashing.” We aim to bridge the gap between regulatory mandates like the EU AI Act and concrete engineering controls, providing an implementation path for verifiable stewardship.
2.2. Evidence Acquisition and Search Strategy
@@
A comprehensive search was conducted across five major academic databases and registers: Scopus ($n=1502$), Web of Science ($n=2798$), IEEE Xplore ($n=279$), ACM Digital Library ($n=656$), and arXiv ($n=78$). To capture the surge in GenAI research following the release of ChatGPT, the search window was limited to publications from November 2022 through December 2025. The search strings utilized an intersection of terms related to Generative AI (e.g., “Large Language Model”, “Foundation Model”, “LLM”) and Responsibility facets (e.g., “AI Safety”, “Trustworthy AI”, “Alignment”, “Hallucination mitigation”). Queries combined RAI concepts (e.g., responsibility, safety, transparency, accountability, fairness, bias mitigation, privacy, equity), with syntax adapted to each database (full strings in Appendix A.1). Syntax was adapted for each database to support field-specific constraints (Title/Abstract/Keywords), and proximity operators were employed where available to enhance precision. To mitigate coverage bias, database queries were complemented with forward and backward snowballing from seed works identified as influential during pilot screening. The review protocol, including sources, screening rules, and the synthesis plan, was prespecified in line with established computer science review guidance (Boland et al., 2017a; Keele et al., 2007)
2.3. Eligibility Criteria
To ensure methodological rigor and relevance to the research questions (RQ1–RQ4), we applied the following criteria:
  • Inclusion Criteria: (1) Peer-reviewed journals, conference proceedings, or high-impact preprints materially shaping RAI discourse; (2) English-language works published within the target 2022–2025 window; and (3) Studies proposing technical frameworks, evaluation metrics, or empirical risk assessments for generative or multimodal systems.
@@ 2.3. Eligibility Criteria
  • Exclusion Criteria: (1) Purely opinionated pieces or editorials without substantive technical argumentation; (2) Short papers or reports lacking sufficient methodological depth; (3) Items whose primary contribution lay outside RAI (e.g., domain applications without safety or governance components); and (4) Redundant duplicates or inaccessible records.
@@
2.4. Study Selection and Screening
As detailed in the PRISMA flow diagram (Fig. 3), the initial search identified 5,313 records. Selection was conducted by three independent reviewers to minimize bias. Following automated deduplication ($n=2125$) and the removal of records marked as ineligible by automation tools ($n=609$) or other technical reasons ($n=53$), 2,562 unique records remained for screening.
Titles and abstracts were screened against the eligibility criteria, resulting in the exclusion of 2,153 items. Subsequently, 372 reports were sought for full-text retrieval and assessed for eligibility. Reports were excluded if they failed to meet quality and technical thresholds ($n=140$) or could not be retrieved ($n=37$). This rigorous multi-reviewer process yielded a final synthesis set of 232 studies. Data extraction followed a thematic pipeline to categorize findings into risks, controls, and evaluation facets as defined by the research objectives.
@@ 3.1. Historical Context
RAI has received a lot of intensified scholarly and policy attention with the rapid application of GenAI (Zirpoli, 2023b; Al-kfairy et al., 2024). GenAI refers to models that synthesize new samples from learned data distributions (van der Zant et al., 2013), spanning language, vision, audio, and multimodal content  (Muneer et al., 2025). Early work combined symbolic rules with probabilistic language models; for example, ELIZA illustrated template driven dialogue in the 1960s (Weizenbaum, 1966). This early era established statistical $n$ gram models (Fink, 2014) and neural sequence modeling for text and speech, laying the groundwork for contemporary large scale generative systems (Qureshi et al., 2025). Figure 4 highlights the most salient milestones for RAI (alignment, evaluation, governance). As shown in the Figure 4. The first major inflection in deep generative modeling arrived with variational autoencoders (VAEs) (Pinheiro Cinelli et al., 2021) and generative adversarial networks (GANs) (Goodfellow et al., 2014), which enabled high-fidelity synthesis and, concurrently, made deepfakes feasible, creating the media integrity challenge that current governance frameworks struggle to address Concurrently, neural sequence-to-sequence models advanced conditional text generation, culminating in the introduction of attention mechanisms and the Transformer (Vaswani et al., 2017), which decoupled sequence modeling from recurrent computation.
A second inflection gave rise to foundation models: large transformers pre-trained in web-scale corpora (e.g., GPT-2/3) and bidirectional encoders (e.g., BERT) that catalyzed transfer across tasks. Alignment advances, including instruction tuning (Peng et al., 2023), reinforcement learning from human feedback (RLHF) (Kaelbling et al., 1996), and constitutional AI (Bai et al., 2022), shifted the emphasis from raw capability to controllability and safety. In vision, diffusion models (Ho et al., 2020) eclipsed GANs for photorealistic generation; when combined with latent compression, they enabled scalable text-to-image synthesis (Rombach et al., 2022).
The most recent phase is multimodal and agentic: vision–language pretraining (e.g., CLIP, Flamingo) (Alayrac et al., 2022), lightweight modality bridging (e.g., BLIP-2) (Li et al., 2023a), and instruction-following multimodal systems (e.g., LLaVA) (Liu et al., 2023b) extend prompting beyond text. Tool use, retrieval-augmented generation (Lewis et al., 2020), and planning/execution patterns operationalize models as agents with external memory and actuation, bringing new safety, privacy, and governance requirements.

### v7

# Who is Responsible? The Data, Models, Users or Regulations? A Comprehensive Survey on Responsible Generative AI for a Sustainable Future

@@ Abstract.
Generative AI is rapidly moving from research to deployment, elevating the need for responsible development, evaluation, and governance. We conduct a PRISMA-guided review of 232 studies (November 2022–December 2025), spanning large language models, vision-language models, diffusion models, and agentic pipelines. We make four contributions: (1) the first survey bridging governance principles, technical evaluation, and domain deployment across all four system types; (2) development of a ten-criterion rubric (C1–C10), scoring major AI safety benchmarks on risk-surface coverage, paired with a policy crosswalk mapping benchmarks to regulatory requirements; (3) establishing twelve lifecycle KPIs, explainability guidance for foundation models, and a catalogue for AI-readye testbed ; and (4) domain-specific analysis across healthcare, finance, education, arts, agriculture, and defense. Three findings emerge: benchmark coverage is dense for bias and toxicity but sparse for privacy, provenance, deepfakes, and system-level failures in agentic settings; evaluations remain largely static and task-local, limiting audit portability; and inconsistent documentation complicates cross-release comparison. We further outline a research agenda, prioritizing adaptive multimodal evaluation, privacy and provenance testing, deepfake risk assessment, calibration reporting, versioned artifacts, and continuous monitoring. This study offers a structured path to align the evaluation of generative AI with the governance needs for a safe and responsible deployment
. Project page: Project page
@@ 1. Introduction
Generative AI (GenAI) systems, including large language models (LLMs), vision language models (VLMs), diffusion models, and emerging agentic pipelines, are rapidly transitioning from research prototypes to production deployments that generate text, multimedia, code, and autonomous actions across healthcare, finance, education, and other sectors (Hajkowicz et al., 2023). This accelerated deployment heightens the need for responsible development practices that address GenAI-specific risks, including hallucination and unverifiability, prompt injection and jailbreak vulnerabilities, data memorization and leakage, deepfake and media integrity threats, and system-level failures in tool-augmented settings.
@@
Responsible AI (RAI) has emerged as a central area of research and practice to address these challenges. RAI frameworks emphasize fairness, bias mitigation, privacy protection, security, and the safeguarding of human rights (Peters et al., 2020). By embedding these principles throughout the AI lifecycle, including design, training, deployment, and governance, RAI seeks to realize the societal benefits of AI while minimizing potential harms. While these principles were initially developed for predictive systems, generative models introduce qualitatively new challenges: outputs are open-ended and difficult to enumerate, stochastic decoding produces non-deterministic behaviour, multimodal synthesis creates cross-modal failure modes, and agentic tool use extends the attack surface beyond the model itself. Despite rapid adoption, with 74% of organizations reporting the use of GenAI, only 26% report having a comprehensive organization-wide responsible AI strategy (Capgemini Research Institute, 2024). This imbalance underscores the need to align the development and deployment of these systems with ethical standards, fairness, privacy, security, and societal values.
@@
There has been substantial progress on RAI from both industry and academia. Nevertheless, notable gaps remain in alignment with human values and in operational practice. The academic literature often separates conceptual frameworks from practical implementations, as summarized in Table 1. Conceptual work provides principles and guidelines, whereas applied studies demonstrate deployment in specific contexts. The recent MIT AI Risk Repository analysis (Slattery et al., 2024) indicates that current AI safety frameworks, including those referenced in Table 4, cover only about 34% of risks identified by governance bodies such as the NIST AI Risk Management Framework (of Standards and (NIST), 2024), the EU AI Act (Outeda, 2024), and ISO/IEC 42001 (ISO/IEC, 2021b). This gap between governance coverage and practical implementation highlights the need for an integrated review that connects frameworks to implementation strategies and to evaluation methods, including explainable AI and AI-ready testbeds.
@@
Figure 1. Full Taxonomy of the Responsible Generative AI landscape.
In this study, we conduct a comprehensive literature review of responsible GenAI (ResGenAI) frameworks, governance guidelines, evaluation benchmarks, and operational risk assessment methodologies. Our scope centres on the application of RAI principles to generative systems, including LLMs, VLMs, diffusion models, and touching on agentic/RAG topics. While we draw on the broader RAI literature for foundational constructs (e.g., fairness, privacy, safety, transparency, accountability), our analysis targets ResGenAI specific risk mechanisms such as hallucination, jailbreaking and tool-use failures and system-level failures in GenAI applications. Despite substantial prior work, ResGenAI guidance remains fragmented across principles, processes, and tools (Table 1). Consequently, there is a pressing need to translate regulatory obligations into concrete engineering controls and to produce portable audit evidence for post-market monitoring. Our key contributions in this work are:
@@
    We present a comprehensive survey that simultaneously spans governance principles, technical evaluation, and domain deployment for GenAI. We synthesize findings across the NIST AI RMF, EU AI Act, ISO/IEC 42001, and other instruments listed in Table 5, covering GenAI topics. As shown in Table 1, no prior survey bridges all four thematic areas (CS theory, application, policy, GenAI) and all four novelty dimensions (taxonomy, scope, evaluation, methods) simultaneously.
@@
    We introduce a multifold-criterion rubric (C1–C10) that scores widely used AI safety benchmarks on risk-surface coverage and assurance quality, and a policy crosswalk (Section 4.2.1) that maps each benchmark to the governance requirements it can plausibly evidence. The closest efforts (Liu et al., 2025a; Ferdaus et al., 2026) partially address evaluation taxonomy or trust assessment, but neither links benchmarks to regulatory audit-evidence needs.
@@
    We apply the ResGenAI lens to real-world applications and domains (Section 5 presents how technical advancements can be effectively aligned with governance frameworks and regulatory requirements.
@@ Methods
(Mehrabi et al., 2021) ACM CSUR 2021 Bias and Fairness x x x x x x
@@
(He et al., 2025) ACM CSUR 2025 LLM Agent Security x x x x
@@
(Liu et al., 2025a) arXiv 2025 LLM Safety Evaluation x x x x
@@
Column definitions CS Theory: conceptual frameworks, taxonomies, or algorithmic contributions. Application: domain-specific deployment or case studies. Policy: governance instruments, regulatory analysis, or compliance guidance. GenAI: explicit focus on generative models (LLMs, VLMs, diffusion, agentic). Taxonomy: structured risk classification beyond a flat listing of concerns. Scope: coverage of multimodal systems and agentic/tool-using pipelines. Evaluation: governance-aligned benchmarking with a scoring rubric. Methods: engineering controls, lifecycle KPIs, or testbed guidance.
The taxonomy presented in Figure 1 provides a structural blueprint for this survey, establishing the thematic connections between high-level technical challenges and domain-specific applications.
@@ 2. Methodology
This review adheres to the PRISMA guidelines (Page et al., 2021) to ensure a transparent, reproducible, and rigorous synthesis of evidence. This study follows established methodological recommendations for performing systematic reviews within the computer science community (Keele and others, 2007). To ensure comprehensive and consistent reporting, the structural framework was modeled after previous extensive surveys in the domain. The mapping procedure utilized in this review is organized into four interconnected phases that covering objectives (Section 2), search strategy (Section 2), eligibility criteria (Section 2), and study selection (Section 2) which are visualized in the methodological flow (Appendix Figure 1) and detailed throughout Section 2.
Research Questions Formulation
@@
This study conducts a literature review of ResGenAI in the post-GPT period and synthesizes the findings. We address the following key questions: RQ1: What are the dominant themes, risks, and controls in ResGenAI and multimodal systems since late 2022? Motivation: The rapid transition of GenAI from research to production has outpaced traditional risk taxonomies, necessitating a contemporary synthesis of how multimodal and agentic systems introduce qualitatively new failure modes beyond predictive ML.
@@
RQ2: Which evaluation facets and metrics are most commonly reported, and where are the gaps? Motivation: Identifying the “Evaluation Gap” is critical for determining if current safety suites adequately model the dynamic pressure applied by real-world adversaries versus merely serving as static, keyword-based refusal scripts.
@@
RQ3: What evidence types (empirical, analytical, normative) underpin claims about RAI effectiveness? Motivation: A structural deficit in evidence portability hinders institutional trust; this question seeks to evaluate if safety claims are grounded in verifiable technical artifacts or subjective, unvalidated appraisals.
@@
RQ4: How can these findings be organized into a practical framework for research and governance? Motivation: Principles without practice lead to “safetywashing.” We aim to bridge the gap between regulatory mandates like the EU AI Act and concrete engineering controls, providing an implementation path for verifiable stewardship.
Evidence Acquisition and Search Strategy
@@
A comprehensive search was conducted across five major academic databases and registers: Scopus ($n=1502$), Web of Science ($n=2798$), IEEE Xplore ($n=279$), ACM Digital Library ($n=656$), and arXiv ($n=78$). To capture the surge in GenAI research following the release of ChatGPT, the search window was limited to publications from November 2022 through December 2025. The search strings utilized an intersection of terms related to GenAI (e.g., “Large Language Model”, “Foundation Model”, “LLM”) and Responsibility facets (e.g., “AI Safety”, “Trustworthy AI”, “Alignment”, “Hallucination mitigation”). Queries combined RAI concepts (e.g., responsibility, safety, transparency, accountability, fairness, bias mitigation, privacy, equity), with syntax adapted to each database (full strings in Appendix 1.1). Syntax was adapted for each database to support field-specific constraints (Title/Abstract/Keywords), and proximity operators were employed where available to enhance precision. To mitigate coverage bias, database queries were complemented with forward and backward snowballing from seed works identified as influential during pilot screening.
Eligibility Criteria
@@
To ensure methodological rigor and relevance to the research questions (RQ1-RQ4), we applied the following criteria: Inclusion Criteria: (1) Peer-reviewed journals, conference proceedings, or high-impact preprints; (2) English-language works published within the target 2022-2025 window; and (3) Studies proposing technical frameworks, evaluation metrics, or empirical risk assessments for GenAI systems. Exclusion Criteria: (1) Purely opinionated pieces or editorials without substantive technical argumentation; (2) Short papers or reports lacking sufficient methodological depth; (3) Items whose primary contribution lay outside RAI (e.g., domain applications without safety or governance components); and (4) Redundant duplicates or inaccessible records.
@@
Study Selection & Screening
As detailed in the PRISMA flow diagram (Fig. 2), the initial search identified 5,313 records. Selection was conducted by three independent reviewers to minimize bias. Following automated deduplication ($n=2125$) and the removal of records marked as ineligible by automation tools ($n=609$) or other technical reasons ($n=53$), 2,526 unique records remained for screening. Titles and abstracts were screened against the eligibility criteria, resulting in the exclusion of 2,117 items. Subsequently, 409 reports were sought for full-text retrieval, of which 37 could not be retrieved. The remaining 372 reports were assessed for eligibility. Reports were excluded if they failed to meet quality and technical thresholds ($n=230$). This rigorous multi-reviewer process yielded a final synthesis set of 142 studies. Data extraction followed a thematic pipeline to categorize findings into risks, controls, and evaluation as defined by the research objectives.
@@ 3.1. Historical Context
RAI has received intensified scholarly and policy attention with the rapid application of GenAI. GenAI refers to models that synthesize new samples from learned data distributions, spanning language, vision, audio, and multimodal content (Muneer et al., 2025). Early work combined symbolic rules with probabilistic language models; ELIZA illustrated template-driven dialogue in the 1960s (Weizenbaum, 1966), while statistical $n$-gram models (Fink, 2014) and neural sequence modeling laid the groundwork for contemporary large-scale generative systems (Qureshi et al., 2025). Figure 3 highlights the most salient milestones for RAI (alignment, evaluation, governance). The first major inflection in deep generative modeling arrived with variational autoencoders (VAEs) (Pinheiro Cinelli et al., 2021) and generative adversarial networks (GANs) (Goodfellow et al., 2014), which enabled high-fidelity synthesis and, concurrently, made deepfakes feasible, creating the media integrity challenge that current governance frameworks struggle to address. Concurrently, neural sequence-to-sequence models advanced conditional text generation, culminating in attention mechanisms and the Transformer (Vaswani et al., 2017), which decoupled sequence modeling from recurrent computation.
A second inflection gave rise to foundation models: large transformers pre-trained on web-scale corpora (e.g., GPT-2/3) and bidirectional encoders (e.g., BERT) that catalyzed transfer across tasks. Alignment advances—instruction tuning (Peng et al., 2023), RLHF (Kaelbling et al., 1996), and constitutional AI (Bai et al., 2022)—shifted emphasis from raw capability to controllability and safety. In vision, diffusion models (Ho et al., 2020) eclipsed GANs for photorealistic generation; combined with latent compression, they enabled scalable text-to-image synthesis (Rombach et al., 2022).
The most recent phase is multimodal and agentic: vision–language pretraining (e.g., CLIP, Flamingo) (Alayrac et al., 2022), lightweight modality bridging (e.g., BLIP-2) (Li et al., 2023b), and instruction-following multimodal systems (e.g., LLaVA) (Liu et al., 2023) extend prompting beyond text. Tool use, retrieval-augmented generation (Lewis et al., 2020), and planning/execution patterns operationalize models as agents with external memory and actuation, introducing new safety, privacy, and governance requirements.